# Universal Acceleration of Quantum Relaxation by Reset-Induced Mpemba Effect

### v2

@@ Abstract
When intrinsic transitions exist ($\Omega\neq 0$), the reset protocol remains effective. In Figure 1 (c) confirms accelerated relaxation for $\Omega=2$, here measured by the L-infinity norm to reduce fluctuations in the curves.
Accelerating relaxation of quantum observables.—In experimental settings, observables rather than full density matrices are typically measured. Our protocol naturally extends to controlling the relaxation dynamics of arbitrary quantum observables. The time evolution of an observable $O$ under unperturbed dynamics reads:
$\langle O\rangle(t)={\rm Tr}[Oe^{\mathcal{L}t}(\rho_{0})]=\langle O\rangle_{ss}+\sum_{k=2}^{d^{2}}c_{k}e^{\lambda_{k}t}{\rm Tr}[OR_{k}].$ (17)
Applying our reset protocol, the modified observable dynamics becomes
$\langle O\rangle^{r}(t)$ $=\langle O\rangle_{ss}+\sum_{k}c_{k}^{r}(t)e^{\lambda_{k}t}\Theta(t_{s}-t){\rm Tr}[OR_{k}]$
$+\sum_{k}c_{k}^{r}(t_{s})e^{\lambda_{k}t}\Theta(t-t_{s}){\rm Tr}[OR_{k}].$ (18)
Thus, the same reset condition $c_{2}^{r}(t_{s})\approx 0$ also ensures fast convergence of $\langle O\rangle(t)$ to its steady-state value $\langle O\rangle_{ss}$. This allows for direct experimental application of our protocol at the level of observables, without requiring full state tomography.
@@
Concluding remarks.—We have introduced a general framework for accelerating relaxation in open quantum systems via quantum reset dynamics, applicable to arbitrary initial states. The proposed protocol is simple to implement: it requires temporarily coupling the system to a reset channel and turning it off after a finite duration $t_{s}$, which can be tuned through the reset rate $r$. Additionally, the protocol can suppress multiple relaxation modes simultaneously, enabling enhanced control over the relaxation dynamics. In the future study, our approach could be extended to certain non-Markovian and periodically driven (Floquet) dissipative systems, and may offer a route to circumvent known trade-offs in relaxation control [18]. Overall, our results establish quantum reset as a versatile and experimentally feasible tool for controlling relaxation timescales in open quantum dynamics.
@@
Appendix
Appendix A Derivation of the modified lindbladian
Acting the time-evolution operator generated by this generalized Lindblad operator on the eigenvalue-expanded form of the initial state
$\rho_{0}=\rho_{ss}+\sum_{k=2}^{d^{2}}\text{Tr}(L_{k}^{\dagger}\rho_{0})R_{k}$ (19)
results in
$\rho^{r}(t)$ $=e^{t\mathcal{L}_{r}}\rho_{0}$
$=e^{t\mathcal{L}_{r}}\rho_{ss}+\sum_{k=2}^{d^{2}}\text{Tr}(L_{k}^{\dagger}\rho_{0})e^{(\lambda_{k}-r)t}R_{k}.$ (20)
To get the detailed form of the first term, note that the stationary state under reset dynamics is given by
$\rho_{ss}^{r}=\lim_{t\rightarrow\infty}\rho^{r}(t)=\rho_{ss}+\sum_{k=2}^{d^{2}}\frac{r\langle\delta|L_{k}^{\dagger}|\delta\rangle}{r-\lambda_{k}}R_{k},$ (21)
and the stationary state has the property that
$e^{t\mathcal{L}_{r}}\rho_{ss}^{r}=\rho_{ss}^{r}.$ (22)
Thus, acting $e^{t\mathcal{L}_{r}}$ on both sides of the Eq.(21) and combining Eq.(22) bring about
$e^{t\mathcal{L}_{r}}\rho_{ss}+\sum_{k=2}^{d^{2}}\frac{r\langle\delta|L_{k}^{\dagger}|\delta\rangle}{r-\lambda_{k}}e^{(\lambda_{k}-r)t}R_{k}$
$=\rho_{ss}+\sum_{k=2}^{d^{2}}\frac{r\langle\delta|L_{k}^{\dagger}|\delta\rangle}{r-\lambda_{k}}R_{k}$ (23)
Appendix B Spectral analysis of the two-state model
The Eq. (13) can be rewritten as a matrix equation for the vector $\vec{\rho}(t)=[\rho_{00}(t),\rho_{01}(t),\rho_{10}(t),\rho_{11}(t)]^{T}$:
$\frac{d\vec{\rho}(t)}{dt}=L\vec{\rho}(t),$ (24)
where $L$ is a $4\times 4$ Liouvillian matrix in the Fock-Liouvillian space, taking the form as
$L=\left(\begin{array}[]{cccc}-\Gamma_{0}&-i\Omega&i\Omega&\Gamma_{1}\\ -i\Omega&-iE-\frac{\Gamma_{0}+\Gamma_{1}}{2}&0&i\Omega\\ i\Omega&0&iE-\frac{\Gamma_{0}+\Gamma_{1}}{2}&-i\Omega\\ \Gamma_{0}&i\Omega&-i\Omega&-\Gamma_{1}\end{array}\right).$ (25)
In the simple case $\Omega=0$, the right eigenvalues and eigenvectors ot the matrix $L$ are given by
$\lambda_{1}$ $=0,\ \vec{R}_{1}=(\Gamma_{1},0,0,\Gamma_{0})^{\text{T}}$
$\lambda_{4}$ $=-\Gamma_{0}-\Gamma_{1},\ \vec{R}_{2}=(-1,0,0,1)^{\text{T}}$
$\lambda_{2}$ $=\frac{1}{2}(-2i-\Gamma_{0}-\Gamma_{1}),\ \vec{R}_{3}=(0,1,0,0)^{\text{T}}$
$\lambda_{3}$ $=\frac{1}{2}(2i-\Gamma_{0}-\Gamma_{1}),\ \vec{R}_{4}=(0,0,1,0)^{\text{T}}$ (26)
Likewise, the left eigenvalues and eigenvectors are given by
$\lambda_{1}^{\star}$ $=0,\ \vec{L}_{1}=(1,0,0,1)^{\text{T}}$
$\lambda_{4}^{\star}$ $=-\Gamma_{0}-\Gamma_{1},\ \vec{L}_{2}=(-\Gamma_{0},0,0,\Gamma_{1})^{\text{T}}$
$\lambda_{2}^{\star}$ $=\frac{1}{2}(2i-\Gamma_{0}-\Gamma_{1}),\ \vec{L}_{3}=(0,1,0,0)^{\text{T}}$
$\lambda_{3}^{\star}$ $=\frac{1}{2}(-2i-\Gamma_{0}-\Gamma_{1}),\ \vec{L}_{4}=(0,0,1,0)^{\text{T}}.$ (27)
Then, by Eq.(8) in the main text, the expression for the critical time $t_{s}$ is given by
$t_{s}=\frac{\ln\left[1+\frac{\Gamma_{1}(r+\Gamma_{0}+\Gamma_{1})\left(\frac{1}{\Gamma_{0}+\Gamma_{0}e^{\beta_{0}E}}-\frac{1}{\Gamma_{0}+\Gamma_{1}}\right)}{r}\right]}{r+\Gamma_{0}+\Gamma_{1}},$ (28)
from which one can see that the critical time wouldn’t be affected by the coherence terms of the initial state.

### v3

@@ Abstract
Example 2.—To illustrate the potential feasibility of our protocol in complex quantum systems, we consider another example, the dissipative transverse-field Ising model (TFIM), a paradigmatic system relevant to various quantum platforms. We numerically study an open dissipative TFIM of length $N$ with Hamiltonian
$H=-J\sum_{i=1}^{N-1}\sigma_{i}^{z}\sigma_{i+1}^{z}-g\sum_{i=1}^{N}\sigma_{i}^{x},$ (18)
and jump operators $J_{i,\downarrow}=\sqrt{\gamma}\,\sigma_{i}^{-},\quad J_{i,\uparrow}=\sqrt{\gamma e^{-\beta}}\;\sigma_{i}^{+}$. Here, $\sigma^{x,z}_{i}$ are Pauli operators, $\sigma_{i}^{-}=|0\rangle\langle 1|$ and $\sigma_{i}^{+}=|1\rangle\langle 0|$. We fix $J=1$, $g=1.2$, $\gamma=0.5$, $\beta=1/k_{B}T=1$ throughout. We choose $\rho_{\delta}=\mathbb{I}/d$ (with $d=2^{N}$). The initial state $\rho_{0}$ is chosen by normalizing $\rho_{\rm ss}+\alpha V_{2}$, with $\rho_{\rm ss}$ being the steady-state and $V_{2}$ being the second normalized eigenvector of the Lindbladian. A larger $\alpha$ implies that $\rho_{0}$ is farther from $\rho_{\rm ss}$. Here, $\rho_{\rm ss}$ is nontrivial, with nonzero entanglement. Parameters: $t_{s}=0.5\,\tau_{2}$ and $t_{s}=0.08\,\tau_{2}$ with $\tau_{2}=1/|\operatorname{Re}\lambda_{2}|$, the relaxation timescale, total duration $T=6\,\tau_{2}$, $\alpha=0.55/0.05$. Notably, $t_{s}$ is chosen arbitrarily, rather than being determined a prior. The accelerated relaxation processes under different reset rate are shown in Fig. 2. We also examine cases with different parameters in [59], where acceleration is consistently observed. In Fig. 2 (b) and (d), genuine Mpemba effects are observed, where the system is first driven away from the steady state by the reset protocol compared to the no-reset case, and then reaches stationarity faster. The existence of such an effect broadens the range of usable reset states.
Some remarks are in order regarding experimental implications. In general, the optimal duration $t_{s}$ of the reset phase cannot be determined a priori from theoretical considerations, as the relevant dynamical details are typically unknown in complex quantum systems. The TFIM example shows how this challenge can be overcome: fine-tuning of $t_{s}$ is not required—any finite value shorter than the native relaxation time leads to substantial speed-up, and $t_{s}$ can be further adjusted a posteriori based on observed dynamics. This robustness, without requiring prior knowledge of the system dynamics and initial states, distinguishes our protocol from previous approaches. Additionally, the reset channel in the example corresponds to a depolarizing channel, which is conceptually straightforward to implement, e.g., by applying an isotropic white-noise field or coupling an infinite-temperature reservoir.
Figure 2: Acceleration of the relaxation of TFIMs. Reset rate $r$ are chosen as $0,1.0,5.0,10.0,20.0$. (a) $N=5$ ($d=32$), $\alpha=0.55$, $t_{s}=0.08\tau_{2}$. (b) $N=5$, $\alpha=0.05$, $t_{s}=0.50\tau_{2}$. (c) $N=6$ ($d=64$), $\alpha=0.55$, $t_{s}=0.08\tau_{2}$. (d) $N=6$, $\alpha=0.05$, $t_{s}=0.50\tau_{2}$. The vertical gray lines indicate $t_{s}$ when the reset channel is removed. Inset: Schematic representation of the dissipative TFIM, with each spin coupled to a thermal bath.
Experimental feasibility in general cases.—The TFIM example illustrates the power of our protocol: the acceleration of relaxation via a depolarizing channel has the potential to be applied to complex quantum platforms. However, in general, it may be necessary to choose other $\rho_{\delta}$ to achieve acceleration. For few-body systems (such as single Rydberg atoms or other multi-level emitters), the protocol may be implemented with quantum reservoir engineering. Such techniques have become increasingly routine for tailoring dissipative dynamics in diverse quantum systems, including atoms [61, 62, 63], trapped ions [48, 64], superconducting circuits [61, 50, 65], and optomechanical setups [66, 67, 68]. For quantum many-body systems, an exact implementation via reservoir engineering generally requires high-order interactions. Such engineered $N$-body dissipators have been put forward theoretically [69, 70] and realized experimentally up to four-body interactions [71].
A complementary route is to approximate the reset channel: combine any available state-preparation protocol for $\rho_{\delta}$, whether reservoir engineering, measurement-based feedback, or other methods, with Trotterization [72, 73, 74] (see End Matter). This strategy sidesteps the need to engineer reset jump operators directly and therefore avoids high-order couplings even in many-body systems. The main experimental challenge then reduces to preparing $\rho_{\delta}$. Encouragingly, fast and high-fidelity state preparation has been reported across diverse platforms [75, 76, 63, 77]. Notably, Ref. [77] realized state preparation in a 35-spin TFIM with Trotterization. This suggests that our protocol may be testable and applicable on existing platforms.
@@
Concluding remarks.—We have introduced a general framework for accelerating relaxation in open quantum systems via temporary reset, applicable to arbitrary initial states. The proposed protocol can suppress multiple relaxation modes simultaneously, enabling enhanced control over the relaxation dynamics. As a practical example, we demonstrated that the relaxation of a TFIM can be significantly accelerated by using a simple depolarizing channel. This example highlights a powerful feature of our protocol: leveraging easily prepared states to accelerate the preparation of states that are difficult to reach. Furthermore, introducing temporary dephasing noise—a different type of channel—can also accelerate relaxation (End Matter). This suggests the broader applicability of our central idea: temporarily coupling the system to various quantum channels may provide a general route to enhanced relaxation. In future studies, our approach could be extended to certain non-Markovian and Floquet dissipative systems (e.g., using ideas in [10, 78]). Overall, our results establish temporary reset as a powerful and experimentally feasible tool for controlling relaxation timescales in open quantum dynamics.
@@
End Matter
Derivation of the modified Lindbladian $\mathcal{L}_{r}$.—Here, we derive Eq. (4) of the main text. Let $\{|\phi_{i}\rangle\}_{i=1}^{d}$ be an orthonormal basis, satisfying $\sum_{i}|\phi_{i}\rangle\!\langle\phi_{i}|=\mathbb{I}$. The reset state is generally written as $\rho_{\delta}=\sum_{\alpha}p_{\alpha}|\psi_{\alpha}\rangle\!\langle\psi_{\alpha}|$. Introduce jump operators
$J_{i,\alpha}^{r}=\sqrt{r\,p_{\alpha}}|\psi_{\alpha}\rangle\!\langle\phi_{i}|,\quad i=1,\dots,d.$
The associated Lindblad term is
$\mathcal{D}_{r}(\rho)=\sum_{i=1}^{d}\!\bigl{(}J_{i}^{r}\rho\,J_{i}^{r\dagger}-\tfrac{1}{2}\{J_{i}^{r\dagger}J_{i}^{r},\rho\}\bigr{)}.$ (19)
Direct evaluation gives
$\sum_{i,\alpha}J_{i,\alpha}^{r}\rho\,J_{i,\alpha}^{r\dagger}=r\,\rho_{\delta}\,\text{Tr}[\rho]$
$-\tfrac{1}{2}\sum_{i,\alpha}\{J_{i,\alpha}^{r\dagger}J_{i,\alpha}^{r},\rho\}=-r\,\rho,$
where $\text{Tr}[\rho]=\sum_{i}\langle\phi_{i}|\rho|\phi_{i}\rangle$, $\sum_{i}|\phi_{i}\rangle\!\langle\phi_{i}|=\mathbb{I}$ and $\langle\psi_{\alpha}|\psi_{\alpha}\rangle=1$ have been used. This yields $\mathcal{D}_{r}(\rho)=r\bigl{[}\text{Tr}[\rho]\,\rho_{\delta}-\rho\bigr{]}$. Hence, we arrive at Eq. (4),
$\mathcal{L}_{r}(\rho)=\mathcal{L}(\rho)+r\text{Tr}[\rho]\,\rho_{\delta}-r\rho.$ (20)
Derivation of the modified coefficients.—Here, we derive Eq. (7), the explicit form of the modified coefficients with reset. Acting the modified semi-group $e^{t\mathcal{L}_{r}}$ on both sides of
$\rho_{0}=\rho_{\rm ss}+\sum_{k=2}^{d^{2}}\text{Tr}(L_{k}^{\dagger}\rho_{0})R_{k}$ (21)
yields
$\rho^{r}(t)=e^{t\mathcal{L}_{r}}\rho_{0}=e^{t\mathcal{L}_{r}}\rho_{\rm ss}+\sum_{k=2}^{d^{2}}\text{Tr}(L_{k}^{\dagger}\rho_{0})e^{(\lambda_{k}-r)t}R_{k}.$ (22)
To proceed, we use an explicit form of the steady state under reset dynamics [45]
$\rho_{\rm ss}^{r}=\lim_{t\rightarrow\infty}\rho^{r}(t)=\rho_{\rm ss}+\sum_{k=2}^{d^{2}}\frac{rd_{k}}{r-\lambda_{k}}R_{k}.$ (23)
Then, expressing the steady-state condition $e^{t\mathcal{L}_{r}}\rho_{\rm ss}^{r}=\rho_{\rm ss}^{r}$ with Eq. (23) and applying Eq. (5), we get
$e^{t\mathcal{L}_{r}}\rho_{\rm ss}+\sum_{k=2}^{d^{2}}\frac{rd_{k}}{r-\lambda_{k}}e^{(\lambda_{k}-r)t}R_{k}=\rho_{\rm ss}+\sum_{k=2}^{d^{2}}\frac{rd_{k}}{r-\lambda_{k}}R_{k},$
from which we solve $e^{t\mathcal{L}_{r}}\rho_{\rm ss}$. Substituting $e^{t\mathcal{L}_{r}}\rho_{\rm ss}$ into Eq. (22) and comparing with Eq. (Accelerating Quantum Relaxation via Temporary Reset: A Mpemba-Inspired Approach), we identify modified coefficients $c_{k}^{r}(t)$ defined in Eq. (7).
Realizing the reset protocol with arbitrary state preparation processes.—We now show that any state preparation protocol could approximately realize the reset channel in our scheme via a Trotterization-based construction. Notably, Trotterization has already been implemented experimentally in superconducting quantum circuits [79], IBM Quantum’s hardware [74] and even quantum many-body platform [77] to generate dissipative dynamics, such as the dissipative TFIM. Moreover, tensor network methods offer promising avenues for extending Trotterization to more complex open quantum systems [72]. We first consider the case where the original dynamics is absent [$\mathcal{L}(\rho)=0$] to gain insight. In this case, the generator corresponding to the reset channel simply reads
$\dot{\rho}(t)=\mathcal{R}[\rho(t)]=r[\rho_{\delta}-\rho(t)].$ (24)
Thus, a direct integration shows that applying the reset channel for a time $t_{s}$ maps an initial state $\rho_{0}$ to a final state $\rho(t_{s})$ according to
$\rho(t_{s})=e^{\mathcal{R}t_{s}}[\rho_{0}]=p\rho_{0}+(1-p)\rho_{\delta},\quad p:=e^{-rt_{s}}.$ (25)
That is, a reset channel with rate $r=-\ln p/t_{s}$ is effectively equivalent to a state preparation protocol for $\rho_{\delta}$, with success probability $1-p$. $p$ is a classical probability generated beforehand. One simply flips a biased coin (or, equivalently, performs any local measurement that yields the desired classical randomness): with probability $1-p$ the system is driven into $\rho_{\delta}$ by any suitable operations and with the probability $p$ it is left untouched. Since there are no restrictions on the preparation method, once $\rho_{\delta}$ is locally preparable (e.g. a separable state), this operation can be implemented purely with local operations, even for many-body systems. Specifically, when $\rho_{\delta}=\mathbb{I}/d$, Eq. (25) is a depolarizing channel, which can be directly realized by the 2-design [80] without flipping a coin.
When $\mathcal{L}\neq 0$, we apply the Lie-Trotter formula [81]
$e^{(\mathcal{L}+\mathcal{R})t}=\lim_{n\to\infty}(e^{\mathcal{R}\frac{t}{n}}e^{\mathcal{L}\frac{t}{n}})^{n}.$ (26)
to realize the protocol. With this formula, the reset protocol during $[0,t_{s}]$ can be approximated as
$e^{(\mathcal{L}+\mathcal{R})t_{s}}[\rho_{0}]\approx(e^{\mathcal{R}\delta t}e^{\mathcal{L}\delta t})^{n}[\rho_{0}],$ (27)
where $\delta t:=t_{s}/n\ll 1$. Using Eq. (25), we have
$e^{\mathcal{R}\delta t}[\rho]\approx(1-p_{s})\rho+p_{s}\rho_{\delta},$ (28)
which can be interpreted as a state preparation mapping from $\rho$ to $\rho_{\delta}$, with success probability $p_{s}:=r\delta t$. Practically, $r$ is determined by $p_{s}$ and $\delta t$, whose values are set by experimentalists.
Thus, the reset protocol may be achieved experimentally by performing the aforementioned state preparation operations stroboscopically: at each discrete time point $t_{i}=i\delta t,\ (i=1,...,n)$ within $[0,t_{s}]$, one applies the state preparation mapping from $\rho(t_{i})$ to $\rho_{\delta}$ with a given probability $p_{s}$. There are no constraints on the preparation time $\tau_{\rm prep}$ for $\rho_{\delta}$, but on average it adds a cost of $np_{s}\tau_{\rm prep}$ to the tailored relaxation timescale. Therefore, it is desirable to choose a $\rho_{\delta}$ that can be prepared efficiently, e.g. the maximally mixed state in our TFIM example. This state is convenient to prepare: coupling to an infinite-temperature bath can achieve it in $\mathcal{O}(\ln N)$ time [82] with minimal resources in a $N$-body system—typically negligible compared to the system’s intrinsic relaxation time. Local control methods could in principle realize it in constant time, though with resources scaling with $N$.
This Trotterization-based method admits a natural interpretation at the level of stochastic trajectories: for a Poissonian reset process with rate $r$, the system is stochastically reset to the target state with probability $r\delta t$, in each small interval $\delta t$.
The approximation error is of order $\mathcal{O}(t_{s}^{2}/n)$ [83, 84, 73]. Explicitly, we establish a rigorous upper bound [59]:
$||e^{(\mathcal{L}+\mathcal{R})t_{s}}-(e^{\mathcal{R}\delta t}e^{\mathcal{L}\delta t})^{n}||\leq\frac{t_{s}^{2}}{2n}||[\mathcal{L,R}]||,$ (29)
where |||||$\cdot$||| | ⋅ | | denotes any norm that is contractive under Lindbladian evolution, such as the trace norm or the diamond norm. This bound holds for arbitrary Lindbladian $\mathcal{L}$ and $\mathcal{R}$. In our specific case,
$[\mathcal{L,R}]\rho:=[\mathcal{LR}-\mathcal{RL}]\rho=r[\mathcal{L}(\rho_{\delta})-\rho_{\delta}].$ (30)
Hence, a large $n$ is not required when $t_{s}\ll 1$, $rt_{s}\ll 1$ or $\rho_{\delta}$ is close to $\rho_{\rm ss}$. A small $t_{s}$ with moderate $r$ offers a practical regime that may yield substantial acceleration and simultaneously reduce the required number of Trotter steps.
Additionally, higher-order approximations could be used to reduce the error for fixed $n$. For instance, the second-order Suzuki-Trotter formula $e^{(\mathcal{L}+\mathcal{R})t_{s}}\approx(e^{\mathcal{L}\delta t/2}e^{\mathcal{R}\delta t}e^{\mathcal{L}\delta t/2})^{n}$ suppresses the error to $\mathcal{O}(t_{s}^{3}/n^{2})$ [83, 73].
Dephasing noise can accelerate the relaxation of the TFIM.—Here, we add dephasing noise along a single axis (typically the $z$-axis) to the TFIM. The corresponding jump operators are $L^{(\phi)}_{i}=\sqrt{\gamma_{\phi}}\sigma_{i}^{z},\quad i=1,\dots,N$. This contributes to the total Lindbladian via an extra term
$\mathcal{L}_{\phi}[\rho]=\gamma_{\phi}\sum_{i=1}^{N}\left(\sigma_{i}^{z}\rho\sigma_{i}^{z}-\rho\right),$ (31)
which could be interpreted as a partial and local reset channel (only the $z$-axis is affected, and only local one-body jump operators are involoved). As shown in Fig. 3, the relaxation is accelerated significantly.
Figure 3: Dephasing noise induce acceleration of relaxation in a 5-site TFIM. Parameters: (a) $J=1.0,\ g=2.0,\ \gamma=0.5,\ \beta=0.1J,\ t_{s}=0.2\tau_{2}$ (b) $J=1.0,\ g=2.0,\ \gamma=0.5,\ \beta=0.1J,\ t_{s}=0.8\tau_{2}$.
A theoretical argument supporting the finding is provided in [59].